# Aligning with Logic: Measuring, Evaluating and Improving Logical Consistency in Large Language Models

### v2

@@ Abstract
Recent research in Large Language Models (LLMs) has shown promising progress related to LLM alignment with human preferences. LLM-empowered decision-making systems are expected to be predictable, reliable and trustworthy, which implies being free from paradoxes or contradictions that could undermine their credibility and validity. However, LLMs still exhibit inconsistent and biased behaviour when making decisions or judgements. In this work, we focus on studying logical consistency of LLMs as a prerequisite for more reliable and trustworthy systems. Logical consistency ensures that decisions are based on a stable and coherent understanding of the problem, reducing the risk of erratic or contradictory outputs. We first propose a universal framework to quantify the logical consistency via three fundamental proxies: transitivity, commutativity and negation invariance. We then evaluate logical consistency, using the defined measures, of a wide range of LLMs, demonstrating that it can serve as a strong proxy for overall robustness. Additionally, we introduce a data refinement and augmentation technique that enhances the logical consistency of LLMs without sacrificing alignment to human preferences. It augments noisy and sparse pairwise-comparison annotations by estimating a partially or totally ordered preference rankings using rank aggregation methods. Finally, we show that logical consistency impacts the performance of LLM-based logic-dependent algorithms, where LLMs serve as logical operators.
@@ 1 Introduction
Recent research in Large Language Models (LLMs; Brown et al. 2020; OpenAI 2023; Anil et al. 2023a; b) has achieved substantial progress concerning their instruction-following capabilities and generating responses aligned with human preferences. Consequently, these advancements have broadened their application to an even wider range of complex natural language tasks across diverse domains, often with minimum or even no supervised data (Kojima et al., 2022). The instruction-following abilities of LLMs are primarily achieved through supervised training with instruction-following data (Wei et al., 2022a; Chung et al., 2022) and reinforcement learning from human feedback (RLHF; Christiano et al. 2017; Stiennon et al. 2020; Ouyang et al. 2022b). During the alignment phases, LLMs are enhanced to align with human values, enabling them to generate responses that better support their decision-making and problem-solving (Dai et al., 2024).
Despite these advancements, key challenges still exist regarding the reliability and trustworthiness of LLMs. Issues such as hallucination (Zhang et al., 2023), bias (Gallegos et al., 2024), and inconsistencies in reasoning (Huang & Chang, 2023) continue to affect their credibility. These limitations hinder the full practical deployment of LLMs, particularly in professional and high-stakes applications where trustworthiness and reliability are crucial. [^1] For instance, in the fields of behavioral economics and psychology systems are traditionally evaluated on two key dimensions: validity and reliability (Schmidt et al., 2000; Guion, 2004; Miller et al., 2023). The foundation of a reliable and trustworthy system is the consistency of its predictions. A consistent system produces explainable and tractable decisions, enhancing its dependability and reliability. In this work, we focus on a key form of consistency in LLMs: logical consistency. This is especially critical for applications requiring structured reasoning and coherent decision-making (Creswell et al., 2023). Logical inconsistencies can lead to unreliable conclusions (Restall, 2002) and even paradoxes (Hyde, 2011), posing significant risks in domains that demand rigorous logical judgment, such as temporal or spatial reasoning (Mostafazadeh et al., 2016b), optimization (Guo et al., 2024), and automated decision systems.
In this paper, we focus on three fundamental aspects of logical consistency: transitivity, commutativity, and negation invariance. These properties are crucial for a coherent understanding of relationships between items, and our goal is to measure, evaluate, and enhance them in LLMs. We propose a universal framework for quantifying these consistency properties, applicable to any number of items, supporting both partial and complete comparisons, and adaptable across various domains. Our empirical investigation spans a diverse set of LLMs, including Llama 3 (Dubey et al., 2024), Gemma (Rivière et al., 2024), and Phi-3 (Abdin et al., 2024), among others. Through extensive evaluations across three representative domains, we find a strong correlation between logical consistency and the internal robustness and reliability of LLMs. This suggests that logical consistency could be a valuable (proxy) indicator of overall model reliability.
To enhance the logical consistency of LLM judgements, we propose a data refinement and augmentation framework that improves noisy and sparse pairwise comparison annotations by estimating preference rankings using rank aggregation methods. The framework then augments the data by ‘logically extrapolating’ additional pairwise comparisons from these rankings. Our results show that LLMs trained with this approach achieve improved internal logical consistency without compromising alignment with human preferences. Furthermore, we demonstrate that, when applied to logic-dependent downstream applications, LLMs with better logical consistency outperform less consistent models. They show better performance and reduced computational requirements when employed as logical operators in a sorting-based preference ranking aggregation algorithm (Liu et al., 2024) that relies on logical coherence.
@@
In sum, our contributions are as follows: 1) We argue that logical consistency, in addition to human alignment, should be taken into account as an essential factor when aligning LLMs’ preferences. We present a general mathematical formulation for quantifying/measuring three key properties of logical consistency: transitivity, commutativity and negation invariance. 2) We then conduct extensive experiments to evaluate logical consistency across a range of state-of-the-art LLMs and analyze the correlations between consistency and models’ reliability. 3) We propose a data refinement and augmentation method for instruction-tuning that enhances logical consistency while maintaining alignment with human preferences and 4) we show that LLMs with improved logical consistency yield improved performance in logic-dependent algorithms where LLMs are employed as logical operators. We will release our source code and refined dataset at ANONYMOUS to facilitate future work.
2 Measuring Logical Consistency
We evaluate the logical consistency of LLMs by assessing their ability to predict logically consistent relations among a set of items. These items could represent diverse entities or events with a uniform relation between them; such a relation might be (i) comparing the preference among response candidates to a query or (ii) determining the causal order of shuffled events, among other possibilities. This evaluation is grounded in relational algebra and order theory (Abiteboul et al., 1995; Imieliński & Lipski Jr, 1984), with the goal of assessing whether the model maintains coherent and self-consistent judgments across its predictions.
To formalize this concept, we define the logical consistency evaluation process by treating an LLM as an operator function that compares pairs of items and outputs a decision on the relation between the items. Let $X=\{x_{1},x_{2},\dots,x_{N}\}$ represent a set of items, and we define a relation (e.g., comparison) function $F:X\times X\rightarrow R$, which compares two items, such as $(x_{i},x_{j})$, and assigns a relational decision $F(x_{i},x_{j})=r$, where $r\in R$ denotes the directional relation between $x_{i}$ and $x_{j}$. For simplicity, we consider $R$ to be a binary relation set, $R=\{r_{ij},r_{ji}\}$, where $r_{ij}$ represents a preferential relation $x_{i}\succ x_{j}$ (i.e., item $x_{i}$ is preferred over item $x_{j}$), and $r_{ji}$ indicates the reverse preference $x_{j}\succ x_{i}$.
In evaluating logical consistency, we focus on whether the function $F$ adheres to the following key properties over the item set $X$: 1) transitivity, 2) commutativity, and 3) negation invariance, as demonstrated in Figure 1. These properties are foundational for logically consistent operations and are critical for determining the reliability and reasoning capability of an LLM (Hamon et al., 2020).
Transitivity ensures that the LLM’s predictions and judgements are internally coherent and do not suffer from logical contradictions within a given context. Commutativity tests whether the model’s decisions are invariant to the order in which items are compared. The model should produce consistent judgments regardless of whether the pair $(x_{i},x_{j})$ or $(x_{j},x_{i})$ is presented first. The negation invariance checks whether the model maintains consistency when dealing with relational negations. Inconsistencies here would indicate a failure to correctly understand and apply the complementary nature of relations. By systematically applying these tests, we are able to assess the extent to which the model’s outputs conform to logically consistent behavior, providing a quantitative proxy measure of its decision reliability.
2.1 Measuring Transitivity
Grounded in our problem setup and definitions above, respecting transitivity implies the following: if a model predicts $A\succ B$ and $B\succ C$, it must also predict $A\succ C$. The ability to make transitive predictions is critical to forming a robust, global understanding of relationships within a set of items. Without transitivity, the model’s reasoning would be prone to contradictions, which could lead to unreliable outcomes in tasks requiring consistent decision-making or ranking (Liu et al., 2024; Li et al., 2019; Qin et al., 2024).
We define that the function $F$ is fully transitive if it does not predict any intransitive relation within the set $X$. This means that if $F(x_{i},x_{j})=r_{ij}$ and $F(x_{j},x_{k})=r_{jk}$, then $F(x_{i},x_{k})=r_{ik}$ must hold for all $i,j,k\in X$. This can be visualized in Figure 3, where we represent the pairwise relations by $F$ as a relation graph. If $F$ is transitive over $X$, the corresponding relation graph should be a Directed Acyclic Graph (DAG). A DAG implies that no cycles exist in the relation graph, denoting that there are no contradictions in the model’s relational judgments. Consequently, to determine if $F$ is fully transitive over the item set $X$, we only have to verify whether the predicted relation graph contains any cycle or not. We show how to construct the relation graph from the judgements of the LLM operator function $F$, and the algorithm to check whether a graph contains cycles in Appendix §A.
@@ 2.1 Measuring Transitivity
We introduce a metric, $s_{tran}(K)$, designed to quantify transitivity over an arbitrary number of items. LLMs often struggle to maintain perfect/full transitivity, especially as the number of items increases. The proposed metric $s_{tran}(K)$ captures the degree of transitivity across subsets of $K$ sampled items, where $3\leq K\leq|X|$. The metric is defined as:
@@
Here, $S_{i}^{K}$ represents a randomly sampled sub-graph of size $K$, and the indicator function $\mathbbm{1}_{acyclic}$ returns $1$ if the sub-graph $S_{i}^{K}$ contains no cycles (i.e., is transitive), and 00 if otherwise. $M$ denotes the total number of the sampled sub-graphs. As the size of the item set increases, the number of possible sub-graph combinations grows exponentially. To manage this complexity, we cap the number of samples to 1,000 sub-graphs to estimate transitivity for larger sets. Therefore, the metric ranges from 00 to $1$, where $1$ represents perfect transitivity.
@@ 2.2 Measuring Commutativity
Commutativity refers to the logical property that ensures the model’s judgments remain consistent when the order of comparison between two items is reversed. Prior studies have shown that LLMs are susceptible to permutation bias, also referred to as positional bias (Wang et al., 2024b; Liusie et al., 2024a). This bias can result in inconsistent outputs when the order of items in a comparison is altered. To measure the degree of commutativity, we propose a metric $s_{comm}$, which evaluates whether the model’s judgment changes when the order of the items is swapped in the prompt. Specifically, it is defined as follows:
$s_{comm}=\frac{2}{|x|\cdot(|x|-1)}\sum_{0<i<j\leq|X|}\mathbbm{1}(F(x_{i},x_{j})=F(x_{j},x_{i})).$ (2)
@@
2.3 Measuring Negation Invariance
Negation invariance tests whether the model maintains consistency when confronted with the negation or inversion of a relational statement. Inconsistencies here could indicate a failure to correctly understand and apply the complementary nature of relations. Previous work suggested that LLMs struggle to automatically infer appropriate inverse relationships when acquiring knowledge (Allen-Zhu & Li, 2023; Berglund et al., 2024). To quantify negation invariance, we propose the metric $s_{neg}$, which examines if the model can correctly reverse its judgement when prompted with a negated relationship between items. The metric is defined as below:
$s_{neg}=\frac{1}{|x|\cdot(|x|-1)}\sum_{0<i,j\leq|X|,i\neq j}\mathbbm{1}(\overset{\neg}{F}(x_{i},x_{j})=\neg F(x_{i},x_{j})).$ (3)
In this formulation, $\neg F(x_{i},x_{j})$ represents the negation of the original relation(e.g., reversing a preference or relational direction). The comparison function $\overset{\neg}{F}(x_{i},x_{j})$ refers to the model’s judgment when explicitly prompted with the negated relation. The indicator function returns $1$ if the model’s response to the negated relation matches the expected negated judgment (i.e., $\overset{\neg}{F}(x_{i},x_{j})=\neg F(x_{i},x_{j})$), and 00 otherwise. We also visualize this comparison in Figure 3. The normalization factor averages across all pairwise permutations in the set $X$, ensuring that $s_{neg}$ ranges from 0 to 1. The maximum score of 1 indicates perfect negation invariance, where the model consistently handles negated relations.
3 Evaluating Logical Consistency of Large Language Models
After defining the measures quantifying the three aspects of logical consistency in Section 2, we now proceed to evaluate LLMs’ judgements from the consistency angle on three representative tasks, each reflecting different levels of subjectivity.
3.1 Evaluation Setup
Tasks and Datasets. We employ three representative tasks to evaluate LLMs’ logical consistency. The first task, abstractive summarization evaluation, uses the SummEval dataset (Fabbri et al., 2021) and focuses on the model’s ability to make preference judgments between summaries, particularly regarding the coherence aspect. The second task, document reranking, leverages the NovelEval dataset (Sun et al., 2023), where LLMs assess the relevance of retrieved documents in response to queries. The final task, temporal event ordering, uses the CaTeRS dataset (Mostafazadeh et al., 2016b) and tests the model’s capability to reason about temporal and causal relationships between events, critical for maintaining consistency in narrative understanding. Further task and dataset details are available in Appendix §B.
Metrics and Reliability Measurement. In Appendix §F, we show the prompt templates used for metrics computation. We compute the logical consistency metrics at the instance level and report averages across the test sets for each task. In addition, we report the human agreement rate (abbreviated as H.) by calculating the pairwise judgement accuracy between judgements made by LLMs and the provided human annotations. It serves as a reference for how closely the model’s judgements align with human judgements. For the measurement of LLMs’ reliability, we perform Monte Carlo Sampling of Chain-of-Thought (Wei et al., 2022b) reasoning outputs, using a temperature of 0.7. Self-agreement is defined as the percentage of outputs that agree with the majority judgment across multiple samples. This measurement ranges from 0.5 to 1, with higher values indicating greater stability in the model’s reasoning processes.
@@ 3.1 Evaluation Setup
Deepseek-chat 74.9 79.4 76.1 82.8 61.5 94.2 90.4 83.0 69.1 81.6 88.2 88.4

### v3

# Aligning with Logic: Measuring, Evaluating and Improving Logical Preference Consistency in Large Language Models

@@ Abstract
Large Language Models (LLMs) are expected to be predictable and trustworthy to support reliable decision-making systems. Yet current LLMs often show inconsistencies in their judgments. In this work, we examine logical preference consistency as a foundational requirement for building more dependable LLM systems, ensuring stable and coherent decision-making while minimizing erratic or contradictory outputs. To quantify the logical preference consistency, we propose a universal evaluation framework based on three fundamental properties: transitivity, commutativity and negation invariance. Through extensive experimentation across diverse LLMs, we demonstrate that these properties serve as strong indicators of judgment robustness. Furthermore, we introduce a data refinement and augmentation technique, REPAIR, that enhances logical consistency while maintaining alignment with human preferences. Finally, we show that improving consistency leads to better performance in LLM-driven logic-based algorithms, reinforcing stability and coherence in decision-making systems.
Machine Learning, ICML
@@ 1 Introduction
Recent research in Large Language Models (LLMs; Brown et al. 2020; OpenAI 2023; Anil et al. 2023a, b) has achieved substantial progress on various tasks, enabling them to generate responses that better support their decision-making and problem-solving (Dai et al., 2024). However, key challenges still exist regarding the reliability and trustworthiness of LLMs. Issues such as hallucination (Zhang et al., 2023), bias (Gallegos et al., 2024), and inconsistencies in reasoning (Huang & Chang, 2023) continue to affect their credibility. These limitations hinder the full deployment of LLMs, particularly in professional and high-stakes applications. [^3] For instance, in the fields of behavioral economics and psychology systems are traditionally evaluated on two key dimensions: validity and reliability (Schmidt et al., 2000; Guion, 2004; Miller et al., 2023).
The foundation of a reliable and trustworthy system is the consistency of its predictions. A consistent system produces explainable and tractable decisions, enhancing its dependability and reliability. In this work, we focus on a key form of consistency in LLMs: logical preference consistency, which is critical for applications requiring structured reasoning and coherent decision-making (Creswell et al., 2023; Hamon et al., 2020). Logical inconsistencies can lead to unreliable conclusions (Restall, 2002) and even paradoxes (Hyde, 2011), posing significant risks in domains that demand rigorous logical judgment, such as temporal or spatial reasoning (Mostafazadeh et al., 2016b), optimization (Guo et al., 2024), and automated decision systems.
@@
This paper examines three key aspects of logical preference consistency in LLMs: transitivity, commutativity, and negation invariance. We propose a universal framework for quantifying these consistency properties, applicable to any number of items, and adaptable across various domains. Our evaluations reveal a strong correlation between consistency and LLMs’ judgment robustness, suggesting that logical preference consistency serves as a useful proxy for preference reliability. To enhance the logical preference consistency of LLMs, we propose REPAIR, a framework that refines noisy pairwise comparisons using rank aggregation and extrapolates additional comparisons logically. Models trained with this approach achieve better internal consistency without sacrificing alignment with human preferences. Moreover, in logic-dependent tasks, these models outperform less consistent ones, demonstrating improved efficiency in sorting-based ranking algorithms (Liu et al., 2024) reliant on logical coherence.
@@
In sum, our contributions are as follows: 1) We highlight the importance of logical preference consistency—alongside human alignment—in aligning LLM preferences. We define mathematical formulations for quantifying/measuring three key consistency properties: transitivity, commutativity and negation invariance. 2) We conduct extensive experiments to evaluate logical preference consistency across state-of-the-art LLMs and analyze its correlation with model reliability. 3) We propose a data refinement and augmentation method for instruction-tuning that enhances logical preference consistency while maintaining alignment with human preferences and 4) we demonstrate that improving logical consistency enhances LLM performance in logic-dependent algorithms where LLMs serve as logical operators. We release our source code and the refined dataset at ANONYMOUS.
@@
2 Measuring Logical Consistency
We evaluate the logical preference consistency of LLMs by assessing their ability to predict logically consistent relations among a set of items. These items could represent diverse entities or events with a uniform relation between them; such a relation might be (i) comparing the preference among response candidates to a query or (ii) determining the causal order of shuffled events, among other possibilities. This evaluation is grounded in relational algebra and order theory (Abiteboul et al., 1995; Imieliński & Lipski Jr, 1984), with the goal of assessing whether the model maintains coherent and self-consistent judgments across its predictions.
To formalize this concept, we define the logical preference consistency evaluation process by treating an LLM as an operator function that compares pairs of items and outputs a decision on the relation between the items. Let $X=\{x_{1},x_{2},\dots,x_{N}\}$ represent a set of items, and we define a relation (e.g., comparison) function $F:X\times X\rightarrow R$, which compares two items, such as $(x_{i},x_{j})$, and assigns a relational decision $F(x_{i},x_{j})=r$, where $r\in R$ denotes the directional relation between $x_{i}$ and $x_{j}$. For simplicity, we consider $R$ to be a binary relation set, $R=\{r_{ij},r_{ji}\}$, where $r_{ij}$ represents a preferential relation $x_{i}\succ x_{j}$ (i.e., item $x_{i}$ is preferred over item $x_{j}$), and $r_{ji}$ indicates the reverse preference $x_{j}\succ x_{i}$.
In evaluating logical consistency, we focus on whether the function $F$ adheres to the following key properties over the item set $X$: 1) transitivity, 2) commutativity, and 3) negation invariance, as demonstrated in Figure 1. Transitivity ensures that the LLM’s predictions and judgements are internally coherent and do not suffer from logical contradictions within a given context. Commutativity tests whether the model’s decisions are invariant to the order in which items are compared. The negation invariance checks whether the model maintains consistent understanding when dealing with relational negations. By systematically applying these tests, we are able to assess the extent to which the model’s judgments conform to logically consistent behavior, providing a quantitative proxy measure of its decision reliability.
2.1 Measuring Transitivity
Grounded in our problem setup and definitions above, transitivity implies that if a model predicts $A\succ B$ and $B\succ C$, it must also predict $A\succ C$. Ensuring transitive predictions is essential for a coherent global understanding of item relationships, preventing contradictions that could undermine reliability in decision-making or ranking tasks (Liu et al., 2024; Li et al., 2019; Qin et al., 2024).
We define that the function $F$ is fully transitive if it does not predict any intransitive relation within the set $X$. This means that if $F(x_{i},x_{j})=r_{ij}$ and $F(x_{j},x_{k})=r_{jk}$, then $F(x_{i},x_{k})=r_{ik}$ must hold for all $i,j,k\in X$. This can be visualized in Figure 3, where we represent the pairwise relations by $F$ as a relation graph. If $F$ is transitive over $X$, the corresponding relation graph should be a Directed Acyclic Graph (DAG).  [^4] A DAG implies that no cycles exist in the relation graph. We assume the comparison $F$ is irreflexive, explained in Appendix §J Consequently, to determine if $F$ is fully transitive over the item set $X$, we only have to verify whether the predicted relation graph contains any cycle. We show how to construct the relation graph from the judgements of the LLM operator function $F$, and the algorithm to check whether a graph contains cycles in Appendix §A.
We introduce $s_{tran}(K)$ to quantify transitivity over an arbitrary number of items. LLMs often struggle to maintain perfect/full transitivity, especially as the number of items increases. The proposed metric $s_{tran}(K)$ captures the degree of transitivity across subsets of $K$ sampled items, where $3\leq K\leq|X|$. The metric is defined as:
@@ 2.1 Measuring Transitivity
Here, $S_{i}^{K}$ represents a randomly sampled sub-graph of size $K$, and the indicator function $\mathbbm{1}_{acyclic}$ returns $1$ if the sub-graph $S_{i}^{K}$ contains no cycles (i.e., is transitive), and 00 if otherwise. $M$ denotes the total number of the sampled sub-graphs. As the size of the item set increases, the number of possible sub-graph combinations grows exponentially. To manage this complexity, we cap the number of samples to 1,000 sub-graphs to estimate transitivity for larger sets. The choice of $M$ is supported by theoretical evidence explained in Appendix §I. Therefore, the metric ranges from 00 to $1$, where $1$ represents nearly perfect transitivity.
@@ 2.2 Measuring Commutativity
Commutativity refers to the logical property that ensures the model’s judgments remain consistent when the order of comparison between two items is reversed. Prior studies have shown that LLMs are susceptible to permutation bias, also referred to as positional bias (Wang et al., 2024b; Liusie et al., 2024a). To measure the degree of commutativity, we propose a metric $s_{comm}$, which evaluates whether the model’s judgment changes when the order of the items is swapped in the prompt. Specifically, it is defined as follows:
$s_{comm}=\frac{2}{|X|(|X|-1)}\sum_{i<j}\mathbbm{1}(F(x_{i},x_{j})=F(x_{j},x_{i})).$ (2)
@@
Deepseek-chat-v2 74.9 79.4 76.1 82.8 61.5 94.2 90.4 83.0 69.1 81.6 88.2 88.4
2.3 Measuring Negation Invariance
Negation invariance tests whether the model maintains consistency when confronted with the negation or inversion of a relational statement. Inconsistencies indicate a failure to correctly understand and apply the complementary nature of relations. Previous work suggested that LLMs struggle to automatically infer appropriate inverse relationships when acquiring knowledge (Allen-Zhu & Li, 2023; Berglund et al., 2024). To quantify negation invariance, we propose the metric $s_{neg}$, which examines if the model can correctly reverse its judgement when prompted with a negated relationship between items. The metric is defined as below:
$s_{neg}=\frac{1}{|X|(|X|-1)}\sum_{\begin{subarray}{c}0<i,j\leq|X|\\ i\neq j\end{subarray}}\mathbbm{1}(\overset{\neg}{F}(x_{i},x_{j})=\neg F(x_{i},x_{j})).$ (3)
In this formulation, $\neg F(x_{i},x_{j})$ represents the negation of the original relation (e.g., reversing a preference or relational direction). $\overset{\neg}{F}(x_{i},x_{j})$ refers to the model’s judgment when explicitly prompted with the negated relation. The indicator function returns $1$ if the model’s response to the negated relation matches the expected negated judgment (i.e., $\overset{\neg}{F}(x_{i},x_{j})=\neg F(x_{i},x_{j})$), and 00 otherwise. We also visualize this comparison in Figure 3. The normalization factor averages across all pairwise permutations in set $X$, ensuring that $s_{neg}$ ranges from 0 to 1. The maximum score of 1 indicates perfect negation invariance, where the model consistently handles negated relations.
3 Evaluating Logical Consistency of LLMs
After defining the measures, we proceed to evaluate LLMs’ judgments from the consistency angle on three representative tasks, each reflecting different levels of subjectivity.
3.1 Evaluation Setup
Tasks and Datasets. We employ three representative tasks to evaluate LLMs’ logical consistency. The first task, abstractive summarization evaluation, uses the SummEval dataset (Fabbri et al., 2021) and focuses on the model’s ability to make preference judgments between summaries, particularly regarding the coherence aspect. The second task, document reranking, leverages the NovelEval dataset (Sun et al., 2023), where LLMs assess the relevance of retrieved documents in response to queries. The final task, temporal event ordering, uses the CaTeRS dataset (Mostafazadeh et al., 2016b) and tests the model’s capability to judge temporal and causal relationships between events, critical for maintaining consistency in narrative understanding. Further task and dataset details are available in Appendix §B.
Metrics and Reliability Measurement. In Appendix §G, we show metric computation details and the prompt templates used. We compute the logical consistency metrics at the instance level and report averages across the test sets for each task. In addition, we report the human agreement rate (abbreviated as H.) by calculating the pairwise judgement accuracy between judgements made by LLMs and the provided human annotations. It serves as a reference for how closely the model’s judgements align with human judgements. For the measurement of LLMs’ reliability, we perform Monte Carlo Sampling of Chain-of-Thought (Wei et al., 2022) reasoning outputs, using a temperature of 0.7. Self-agreement is defined as the percentage of outputs that agree with the majority judgment across multiple samples. This measurement ranges from 0.5 to 1, with higher values indicating greater stability in the model’s judgments.